# Non-stabilizerness in open XXZ spin chains: Universal scaling and dynamics

### v1

@@ Abstract
Magic, or non-stabilizerness, is a crucial quantum resource, yet its dynamics in open quantum systems remain largely unexplored. We investigate magic in the open XXZ spin chain under either boundary gain and loss, or bulk dephasing using the stabilizer Rényi entropy $M_{2}$. To enable scalable simulations of large systems, we develop a novel, highly efficient algorithm for computing $M_{2}$ within the matrix product states formalism while maintaining constant bond dimension—an advancement over existing methods. For boundary driving, we uncover universal scaling laws, $M_{2}(t)\sim t^{1/z}$, linked to the dynamical exponent $z$ for several distinct universality classes. We also disentangle classical and quantum contributions to magic by introducing a mean-field approximation for magic, thus emphasizing the prominent role of quantum critical fluctuations in non-stabilizerness. For bulk dephasing, dissipation can transiently enhance magic before suppressing it, and drive it to a nontrivial steady-state value. These findings position magic as a powerful diagnostic tool for probing universality and dynamics in open quantum systems.
Introduction.
Non-stabilizerness quantifies the deviation of a quantum state from the set of stabilizer states, serving as a key resource for quantum advantage [1, 2, 3, 4] in tasks such as quantum computing and simulation [5, 6, 7, 8]. In closed systems, unitary evolution generates and propagates magic through coherent interference, with chaotic dynamics typically leading to faster magic growth compared to localized phases [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. However, in open quantum systems, the interplay between coherent dynamics and decoherence fundamentally alters the behavior of magic. While dissipation and dephasing generally suppress magic by destroying quantum coherences, driven-dissipative systems can exhibit non-equilibrium steady states (NESS) that preserve or even generate magic, particularly in the presence of interactions [19, 20, 21]. Understanding the dynamics of magic in open systems is therefore crucial for harnessing quantum resources in realistic, noisy environments [22, 23, 24].
@@ Introduction.
First, we develop a highly efficient algorithm for computing the stabilizer Rényi entropy (magic) in open quantum systems, which maintains a constant bond dimension throughout the calculation. This represents a significant improvement over existing methods, as it avoids the bond dimension growth typically associated with tensor network approaches [32, 33, 34].
For boundary gain and loss, we identify universal scaling law for magic in the XXZ chain, showing $M_{2}(t)\sim t^{1/z}$ with $z$ the dynamical critical exponent. It reflects various transport regimes, including ballistic $(z=1)$, diffusive $(z=2)$, and even the Kardar-Parisi-Zhang (KPZ) dynamics [35, 36, 37] at the isotropic point with $z=3/2$. To disentangle classical and quantum contributions to non-stabilizerness, we introduce a mean-field approximation for magic that eliminates quantum correlations while preserving local magnetization contributions. This approach provides a classical baseline for magic, allowing us to quantify the role of quantum correlations, and identify their impact on transport dynamics and magic generation.
In the presence of bulk dephasing when all sites are coupled to environment through coupling $\gamma_{z}$, we find purely dissipative dynamics with no coherent evolution at the spin rotational invariant limit. This occurs when starting from a fully polarized state in any direction. In general, we find that the time evolution is governed by two distinct timescales: the dephasing timescale $\tau_{d}\sim 1/\gamma_{z}$, which drives the exponential suppression of magic, and the coherence timescale $\tau_{c}$. The latter is the inverse of the SU(2) symmetry breaking interaction and controls transient oscillations in magic. Remarkably, dephasing can transiently enhance magic before driving it to zero. However, in magnetization-conserving systems, magic does not fully decay but stabilizes at a non-trivial steady-state value. Furthermore, when the dynamics is restricted to the zero magnetization sector, starting from certain initial states such as the Néel configuration, magic exhibits a robust power-law decay over time, reflecting the underlying gapless nature of the Lindbladian [38, 39].
Stabilizer Rényi entropy in open systems.
Figure 1: (a) Illustration of the MPS structure for the vectorized density matrix, with explicit site labels shown. (b) Depiction of the Pauli vector construction and its corresponding MPS structure. The red dots represent the Pauli tensors. (c) After contracting the physical indices, the Pauli vector is expressed as an MPS with the same bond dimension $\chi$ as the original vectorized density matrix.
@@
where $P$ represents elements of the $N$-qubit Pauli group [40, 41, 42], and the Pauli coefficients $c_{P}$ are $c_{P}=\mathrm{Tr}(\rho P)$. In open systems, the Pauli spectrum exhibits both delocalization due to coherent dynamics and decay from non-unitary processes. Even for closed systems, calculating the full Pauli spectrum is computationally expensive for large $N$, due to the exponential growth of the Pauli group. Efficient methods like Monte Carlo sampling [43] and tensor networks [44, 32, 45] allow for practical evaluation of $M_{2}$ in closed systems.
@@
We introduce a new approach to compute magic in open systems, using a vectorized density matrix structure while maintaining a constant bond dimension. This method, extending the framework of Ref. [45], avoids the bond dimension increase required by closed-system algorithms, offering a scalable and computationally efficient tool for quantifying magic in open quantum systems.
MPS in the vectorized basis.
@@
Figure 2: (a,b,c) Time evolution of magic for different anisotropy parameters $\Delta$ and for various system sizes. The insets in each panel present the light cone formation in magnetization for the respective $\Delta$ as a function of position $x=2j/L-1$, with $j$ numbering the sites. The initial density matrix corresponds to the infinite-temperature state, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$. For $\Delta=0$, the magic and magnetization light cone exhibit a linear growth, characteristic for the ballistic transport regime. At $\Delta=1$, the growth follows a $t^{2/3}$ scaling, indicative of the KPZ universality class, while for $\Delta=2$, it becomes diffusive with a dynamical exponent $z=2$. The legend and colorbar are shared for all panels.
@@
Boundary driven XXZ spin chain.
where $X_{j},Y_{j}$, and $Z_{j}$ are Pauli operators at site $j$, and $\Delta$ is the anisotropy parameter. We investigate the generation and depletion of magic in the XXZ chain coupled to external reservoirs. Specifically, we focus on boundary-driven dynamics, where interactions with reservoirs at the chain’s edges drive the system into a nonequilibrium steady state (NESS) [25, 49] as displayed in Fig. 3(a). The dissipation at the boundaries is described by the Lindblad operators
Following the quench, the system undergoes a dynamical evolution, eventually reaching a nonequilibrium steady state (NESS) characterized by a constant magnetization current [49]. The emergence of two expanding light cones, originating from the system’s boundaries, is clearly visible in the magnetization profiles of Fig. 2(a,b,c), where their spatial spread follows the dynamical scaling $x\sim t^{1/z}$. The exponent $z$ dictates the transport behavior, distinguishing different universality classes. A crucial aspect of this evolution is the initial state choice. The system is initialized in the infinite-temperature density matrix, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$, which ensures an uncorrelated and homogeneous starting point. This setup allows the observed transport properties and the associated dynamical scaling to emerge purely from the interplay between unitary evolution and boundary driving, rather than from any pre-existing correlations. In particular, it guarantees that the KPZ universality class at $\Delta=1$ arises intrinsically from the system’s dynamics, reinforcing the robustness of the observed scaling behavior.
For $\Delta<1$, the system exhibits ballistic transport [50], with a linear light cone ($z=1$). At the critical point $\Delta=1$, the dynamics enters the Kardar-Parisi-Zhang (KPZ) universality class [37, 51, 52], where transport becomes superdiffusive with a characteristic light cone scaling $z=3/2$. This anomalous behavior reflects the emergence of strong correlations and fluctuations. In contrast, for $\Delta>1$, transport is diffusive ($z=2$), resulting in a broader and slower-growing light cone.
Remarkably, the time evolution of magic follows the same universal scaling as the magnetization profile, with $M_{2}(t)\sim t^{1/z}$. Rescaling by $\mu^{2}$ and by the system size $L$ reveals a universal collapse of the magic dynamics across different transport regimes, underscoring its role as a robust measure of the system’s universality class. In the KPZ regime ($\Delta=1$), magic exhibits a characteristic $t^{2/3}$ growth, in precise agreement with the expected scaling. This highlights magic as a sensitive probe of quantum transport, capturing the emergence of quantum correlations that define the KPZ universality class, beyond what is accessible through conventional observables like magnetization currents. Similar to the magnetization profile, magic does not develop uniformly but instead emerges within the two light cones propagating from the system’s edges. These expanding regions serve as the primary sources for magic generation before it eventually saturates in the NESS. As shown in Fig. 2 (insets), the transition to the steady state coincides with the collision of the two light cones at the center, marking the completion of the transport-driven magic buildup.
One of the key findings of this work is the emergence of a universal scaling law for magic, as demonstrated in Fig. 2 (d,e,f). Our results establish that the magic follows the scaling form
$M_{2}(L,t)\sim L\,f\left(\frac{t}{L^{z}}\right),$ (10)
where the scaling function $f(x)$ governs the dynamical evolution across different regimes. In the early-time regime ($t\ll L^{z}$), magic exhibits a universal growth $M_{2}(t)\sim t^{1/z}$, directly reflecting the system’s transport properties, $f(x\ll 1)\sim x^{1/z}$. At late times ($t\gg L^{z}$), magic saturates as $M_{2}\sim L$, signaling the approach to the nonequilibrium steady state, which gives $f(x\gg 1)\sim const$ for the scaling function. This universal scaling structure provides a unifying framework for understanding its growth across different dynamical universality classes.
Figure 4: (a) Sketch of the bulk dephasing modeled with sites coupled to an external reservoir via the jump operator $S^{z}$. (b, c) Time evolution under dephasing affecting all qubits at the isotropic point $\Delta=1$. (b) Starting from an initial density matrix with zero magic, $\rho(0)=(|+\rangle\langle+|)^{\otimes L}$, the magic initially increases before undergoing an exponential decay to zero at late times. The inset displays the magic density in the NESS state $M_{2}(t_{\infty},\theta)$ as a function of the initial polarization angle $\theta$. (c) When the initial density matrix is $\rho(0)=(|T\rangle\langle T|)^{\otimes L}$, the system possesses a high degree of magic. Due to conservation of the total $S_{z}$ spin component there is a non-zero magic in the NESS $M_{2}(t_{\infty})=\log_{2}(6/5)$, subtracted in $(b)$ from $M_{2}(t)$. The dashed line denotes the analytical result from Eq. (11).
Mean-field approximation for magic.
Figure 3 (b) highlights the fundamental role of transport in determining the steady-state magic. In the ballistic regime ($\Delta\ll 1$), quantum correlations are dominant, maximizing the gap between $M_{2}(t_{\infty})$ and $M_{2}^{\rm{MF}}(t_{\infty})$. Since the magnetization profile remains nearly unchanged for $\Delta\leq 0.6$ [see Fig. 3 (c)], the mean-field magic stays almost constant. The small amplitude of $M_{2}^{\rm{MF}}(t_{\infty})$ reflects the local magnetization, which almost vanishes in the bulk, leaving boundary sites as the primary contributors. Near $\Delta\approx 1$, both measures of magic rise sharply, signaling enhanced entanglement and a fundamental shift in the magnetization profile. In the diffusive regime, the gap between full and mean-field magic narrows, reflecting the suppression of quantum correlations.
The explicit decomposition of magic into classical and quantum components is a key achievement of this work, as it provides a direct measure of non-stabilizerness due to quantum correlations. Notably, despite being entirely classical, $M_{2}^{\rm{MF}}(t)$ exhibits the same scaling behavior as the full magic [39], confirming that universal transport signatures persist even in the absence of quantum correlations.
Quantum Dephasing.
Heisenberg chain.
When rescaled by system size, magic dynamics collapse onto a universal curve, matching the single-qubit behavior. The inset of Fig. 4(a) shows a polar plot of the steady-state magic.
Conclusions.
For boundary driving, we reveal a universal scaling law for magic, $M_{2}(t)\sim t^{1/z}$, directly tied to the dynamical exponent $z$ of the underlying transport regime, and demonstrate that magic serves as a sensitive probe of quantum transport, including the Kardar-Parisi-Zhang universality class.
We introduce a magic mean-field approach, which isolates the classical contribution to magic through local expectation values. The difference between full and magic mean-field serves as a direct probe of non-stabilizerness generated by the many-body correlations built into the system.
For bulk dephasing, we demonstrate that dissipation can transiently enhance magic before ultimately driving it toward its steady state value. Magic exhibits an initial growth, peaks around a characteristic dephasing timescale $t\approx 1/\gamma_{z}$, and then decays exponentially. In systems with conserved magnetization, however, magic is not fully erased and stabilizes at a non-trivial steady-state value. In the Heisenberg limit, we derive an analytical expression capturing this time dependence. Away from the SU(2) point, the dynamics are governed by two key timescales: the dephasing time $\tau_{d}\sim 1/\gamma_{z}$, which controls the exponential decay, and the coherence time $\tau_{c}\sim 1/J|\Delta-1|$, which accounts for transient oscillations away from the $SU(2)$ point. Intriguingly, when the system is initialized in e.g. the Néel or $(|+\rangle\langle+|)^{\otimes L}$ configurations within the zero-magnetization sector—magic decays as a power law in time, revealing signatures of the gapless nature of the Lindbladian.
Acknowledgements.

### v2

# Nonstabilizerness in open XXZ spin chains: Universal scaling and dynamics

@@ Abstract
Magic, or nonstabilizerness, is a crucial quantum resource, yet its dynamics in open quantum systems remains largely unexplored. We investigate magic in the open XXZ spin chain under either boundary gain and loss, or bulk dephasing using the stabilizer Rényi entropy $M_{2}$. To enable scalable simulations of large systems, we develop a novel, highly efficient algorithm for computing $M_{2}$ within the matrix product state formalism while maintaining constant bond dimension—an advancement over existing methods. For boundary driving, we uncover universal scaling laws, $M_{2}(t)\sim t^{1/z}$, linked to the dynamical exponent $z$ for several distinct universality classes. We also disentangle classical and quantum contributions to magic by introducing a mean-field approximation for magic, thus emphasizing the prominent role of quantum-critical fluctuations in nonstabilizerness. For bulk dephasing, dissipation can transiently enhance magic before suppressing it and drive it to a nontrivial steady-state value. These findings position magic as a powerful diagnostic tool for probing universality and dynamics in open quantum systems.
I Introduction
Nonstabilizerness quantifies the deviation of a quantum state from the set of stabilizer states, serving as a key resource for quantum advantage [1, 2, 3, 4] in tasks such as quantum computing and simulation [5, 6, 7, 8]. In closed systems, unitary evolution generates and propagates magic through coherent interference, with chaotic dynamics typically leading to faster magic growth compared to localized phases [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However, in open quantum systems, the interplay between coherent dynamics and decoherence fundamentally alters the behavior of magic. While dissipation and dephasing generally suppress magic by destroying quantum coherences, driven-dissipative systems can exhibit nonequilibrium steady states (NESS) that preserve or even generate magic, particularly in the presence of interactions [20, 21, 22]. Understanding the dynamics of magic in open systems is, therefore, crucial for harnessing quantum resources in realistic, noisy environments [23, 24, 25].
@@ I Introduction
First, we develop a highly efficient algorithm for computing the stabilizer Rényi entropy (magic) in open quantum systems. While previous approaches scale as $M^{4}$ or $M^{2}$ in the bond dimension $M$ [33, 34, 35], our method keeps the bond dimension constant, resulting in linear memory scaling with $M$ and significantly reduced computational cost. This ensures that large bond dimensions, up to $M\sim 1000$, can be handled efficiently in practice.
With this algorithm in hand, we next investigate the open XXZ spin chain. For boundary gain and loss, we identify a universal scaling law for magic in the XXZ chain, showing $M_{2}(t)\sim t^{1/z}$ with $z$, being the dynamical critical exponent. It reflects various transport regimes, including ballistic $(z=1)$, diffusive $(z=2)$, and even the Kardar-Parisi-Zhang (KPZ) dynamics [36, 37, 38] at the isotropic point with $z=3/2$. Having established the boundary-driven case, we then move to bulk dephasing, which reveals qualitatively different physics and additional insights. To disentangle classical and quantum contributions to nonstabilizerness, we introduce a mean-field approximation for magic that eliminates quantum correlations while preserving local magnetization contributions. This approach provides a classical baseline for magic, allowing us to quantify the role of quantum correlations and identify their impact on transport dynamics and magic generation.
In the presence of bulk dephasing when all sites are coupled to environment through coupling $\gamma_{z}$, we find purely dissipative dynamics with no coherent evolution at the spin rotational invariant limit. This occurs when starting from a fully polarized state in any direction. In general, we find that the time evolution is governed by two distinct timescales: the dephasing timescale $\tau_{d}\sim 1/\gamma_{z}$, which drives the exponential suppression of magic, and the coherence timescale $\tau_{c}$. The latter is the inverse of the SUSU(2) symmetry-breaking interaction and controls transient oscillations in magic. Remarkably, dephasing can transiently enhance magic before driving it to zero. However, in magnetization-conserving systems, magic does not fully decay but stabilizes at a nontrivial steady-state value. Furthermore, when the dynamics is restricted to the zero magnetization sector, starting from certain initial states such as the Néel configuration, magic exhibits a robust power-law decay over time, reflecting the underlying gapless nature of the Lindbladian [39, 40].
II Stabilizer Rényi entropy in open systems
Figure 1: (a) Illustration of the MPS structure for the vectorized density matrix, with explicit site labels shown. (b) Depiction of the Pauli vector construction and its corresponding MPS structure. Red dots represent the Pauli tensors. (c) After contracting the physical indices, the Pauli vector is expressed as an MPS with the same bond dimension $\chi$ as the original vectorized density matrix.
@@ II Stabilizer Rényi entropy in open systems
where $P$ represents elements of the $N$-qubit Pauli group [41, 42, 43] and the Pauli coefficients $c_{P}$ are $c_{P}=\mathrm{Tr}(\rho P)$. In open systems, the Pauli spectrum exhibits both delocalization due to coherent dynamics and decay from nonunitary processes. Even for closed systems, calculating the full Pauli spectrum is computationally expensive for large $N$, due to the exponential growth of the Pauli group. Efficient methods like Monte Carlo sampling [44] and tensor networks [45, 33, 46] allow for practical evaluation of $M_{2}$ in closed systems.
@@
We introduce a new approach to compute magic in open systems, using a vectorized density matrix structure while maintaining a constant bond dimension. Details of the algorithm are presented in the Appendix.
This method, extending the framework of Ref. [46], avoids the bond dimension increase required by closed-system algorithms, offering a scalable and computationally efficient tool for quantifying magic in open quantum systems.
@@
Figure 2: (a,b,c) Time evolution of magic for different anisotropy parameters $\Delta$ and for various system sizes. The insets in each panel present the light-cone formation in magnetization for the respective $\Delta$ as a function of position x=2j/L1x=2j/L-1, with $j$ numbering the sites. The dashed lines in the insets mark the boundaries of the emerging light cones, defined as the points where the magnetization reaches $1/4$ of its maximum value. The initial density matrix corresponds to the infinite-temperature state, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$. For $\Delta=0$, the magic and magnetization light cone exhibit a linear growth, characteristic for the ballistic transport regime. At $\Delta=1$, the growth follows a $t^{2/3}$ scaling, indicative of the KPZ universality class, while for $\Delta=2$, it becomes diffusive with a dynamical exponent $z=2$. The legend and color bar are shared for all panels.
We represent the vectorized density matrix $|\rho\rangle\!\rangle$ as an MPS [47, 48, 49], with explicit tracking of the physical index pairs $(s_{i},s^{\prime}_{i})$:
$|\rho\rangle\!\rangle=\sum_{s_{i},s^{\prime}_{i}}A_{1}^{s_{1},s^{\prime}_{1}}A_{2}^{s_{2},s^{\prime}_{2}}\cdots A_{L}^{s_{L},s^{\prime}_{L}}|s_{1}s^{\prime}_{1}\dots s_{L}s^{\prime}_{L}\rangle.$ (2)
This form allows for efficient manipulation while preserving the system’s structure and normalization properties. To compute the stabilizer Rényi entropy, we construct a matrix product operator (MPO) $W$ that encodes the squared action of local Pauli operators according to Eq. (13). The entropy is then given by
Here, $|P\rho\rangle$ is a Pauli-transformed version of $|\rho\rangle\!\rangle$, constructed such that the bond dimension remains unchanged. The structures of $|\rho\rangle\!\rangle$ and $|P\rho\rangle$ are illustrated in Fig. 1.
III Boundary-driven XXZ spin chain
where $X_{j},Y_{j}$, and $Z_{j}$ are Pauli operators at site $j$, and $\Delta$ is the anisotropy parameter. We investigate the generation and depletion of magic in the XXZ chain coupled to external reservoirs. Specifically, we focus on boundary-driven dynamics, where interactions with reservoirs at the chain’s edges drive the system into a nonequilibrium steady state [26, 50] as displayed in Fig. 3(a). The dissipation at the boundaries is described by the Lindblad operators
Following the quench, the system evolves toward a NESS characterized by a steady magnetization current [50]. During this evolution, two distinct light-cone structures emerge from the system’s boundaries and propagate into the bulk. These expanding fronts are clearly visible in the magnetization profiles shown in the insets of Figs. 2 (a)-2(c). Their spatial extent $\Delta x$ exhibits a universal scaling $\Delta x\sim t^{1/z}$, reflecting the underlying transport regime: ballistic ($z=1$) in the easy-plane phase, KPZ-type scaling ($z=3/2$) at the isotropic point, and diffusive spreading ($z=2$) in the easy-axis regime. The exponent $z$ dictates the transport behavior, distinguishing different universality classes. A crucial aspect of this evolution is the initial state choice. The system is initialized in the infinite-temperature density matrix, $|\rho_{0}\rangle\!\rangle=|\mathbb{1}_{\infty}\rangle\!\rangle$, which ensures an uncorrelated and homogeneous starting point. This setup allows the observed transport properties and the associated dynamical scaling to emerge purely from the interplay between unitary evolution and boundary driving, rather than from any pre-existing correlations. In particular, it guarantees that the KPZ universality class at $\Delta=1$ arises intrinsically from the system’s dynamics, reinforcing the robustness of the observed scaling behavior.
For $\Delta<1$, the system exhibits ballistic transport [51], with a linear light cone ($z=1$). At the critical point $\Delta=1$, the dynamics enters the Kardar-Parisi-Zhang universality class [38, 52, 53], where transport becomes superdiffusive with a characteristic light-cone scaling $z=3/2$. This anomalous behavior reflects the emergence of strong correlations and fluctuations. In contrast, for $\Delta>1$, transport is diffusive ($z=2$), resulting in a broader and slower-growing light cone.
Remarkably, the time evolution of magic follows the same universal scaling as the magnetization profile, with $M_{2}(t)\sim t^{1/z}$. Rescaling by $\mu^{2}$ and the system size $L$ reveals a universal collapse of the magic dynamics across different transport regimes, underscoring its role as a robust measure of the system’s universality class. In the KPZ regime ($\Delta=1$), magic exhibits a characteristic $t^{2/3}$ growth, in precise agreement with the expected scaling. This highlights magic as a sensitive probe of quantum transport, capturing the emergence of quantum correlations that define the KPZ universality class, beyond what is accessible through conventional observables like magnetization currents. Similar to the magnetization profile, magic does not develop uniformly but instead emerges within the two light cones propagating from the system’s edges. These expanding regions serve as the primary sources of magic generation before it eventually saturates in the NESS. As shown in Fig. 2 (insets), the transition to the steady state coincides with the collision of the two light cones at the center, marking the completion of the transport-driven magic buildup.
One of the key findings of this work is the emergence of a universal scaling law for magic, as demonstrated in Fig. 2 (a,b,c). Our results establish that the magic follows the scaling form
$M_{2}(L,t)\sim L\,f\left(\frac{t}{L^{z}}\right),$ (6)
where the scaling function $f(x)$ governs the dynamical evolution across different regimes. In the early-time regime ($t\ll L^{z}$), magic exhibits a universal growth $M_{2}(t)\sim t^{1/z}$, directly reflecting the system’s transport properties, $f(x\ll 1)\sim x^{1/z}$. At late times ($t\gg L^{z}$), magic saturates as $M_{2}\sim L$, signaling the approach to the nonequilibrium steady state, which gives $f(x\gg 1)\sim\text{const}$ for the scaling function. The saturation time $t_{0}$, at which $M_{2}$ reaches a plateau, is closely connected to the growth of correlations in the system; it marks the point where the correlation length, constrained by the system size, approaches $\xi\sim L$. This universal scaling structure provides a unifying framework for understanding its growth across different dynamical universality classes.
Figure 4: (a) Sketch of the bulk dephasing modeled with sites coupled to an external reservoir via the jump operator $S^{z}$. (b, c) Time evolution under dephasing affecting all qubits at the isotropic point $\Delta=1$. (b) Starting from an initial density matrix with zero magic, $\rho(0)=(|+\rangle\langle+|)^{\otimes L}$, the magic initially increases before undergoing an exponential decay to zero at late times. The inset displays the magic density in the NESS state $M_{2}(t_{\infty},\theta)$ as a function of the initial polarization angle $\theta$. (c) When the initial density matrix is $\rho(0)=(|T\rangle\langle T|)^{\otimes L}$, the system possesses a high degree of magic. Due to conservation of the total $S_{z}$ spin component there is a nonzero magic in the NESS $M_{2}(t_{\infty})=\log_{2}(6/5)$, subtracted in (c) from $M_{2}(t)$. The dashed line denotes the analytical result from Eq. (7). (d, e) Power-law decay of magic $M_{2}$ for various initial states within the $\langle S_{z}\rangle=0$ sector: (d) a product state $\rho(0)=(|+\rangle\langle+|)^{\otimes L}$ and (e) a Néel state, both evaluated for $\gamma_{z}=1$ and $L=96$. At the isotropic point $\Delta=1$, the data in (d) recovers the exponential decay already presented in panel (b).
To disentangle classical from quantum contributions, we further introduce a mean-field approximation, which provides a natural baseline for comparison.
IV Mean-field approximation for magic
Figure 3(b) highlights the fundamental role of transport in determining the steady-state magic. In the ballistic regime ($\Delta\ll 1$), quantum correlations are dominant, maximizing the gap between $M_{2}(t_{\infty})$ and $M_{2}^{\rm{MF}}(t_{\infty})$.
For small values of $\Delta$ (specifically $\Delta\lesssim 0.6$), the magnetization profile remains largely unchanged [as presented in Ref. [40], Fig. 6(b); see also Ref. [54]], resulting in an almost constant value of the mean-field magic. In this regime, the local magnetization is negligible throughout the bulk and only remains significant near the boundaries. Consequently, the mean-field magic $M_{2}^{\rm{MF}}(t_{\infty})$, which depends solely on local expectation values, is primarily supported by contributions from the edge sites, leading to its overall smaller magnitude, in contrast to the total $M_{2}$. Near $\Delta\approx 1$, both measures of magic rise sharply, signaling enhanced entanglement and a fundamental shift in the magnetization profile. In the diffusive regime, the gap between full and mean-field magic narrows, reflecting the suppression of quantum correlations. In the large-$\Delta$ limit, the steady-state magnetization profile converges to $\langle Z(x)\rangle\sim\arcsin(x)$, with vanishing transverse components [54]. While the mean-field magic is computed only from local observables, the full quantum magic retains contributions from nonlocal correlations such as $\langle\cdots Z_{i}\cdots Z_{j}\cdots\rangle$, resulting in a finite and persistent gap between the full and mean-field values even as $\Delta\to\infty$, as confirmed by our simulations for large $\Delta$.
The explicit decomposition of magic into classical and quantum components is a key achievement of this work, as it provides a direct measure of nonstabilizerness due to quantum correlations. Notably, despite being entirely classical, $M_{2}^{\rm{MF}}(t)$ exhibits the same scaling behavior as the full magic [40], confirming that universal transport signatures persist even in the absence of quantum correlations.
V Quantum Dephasing
VI Heisenberg chain
When rescaled by system size, magic dynamics collapse onto a universal curve, matching the single-qubit behavior. The inset of Fig. 4(a) shows a polar plot of the steady-state magic. Taken together, these complementary perspectives build a coherent picture of how magic behaves in open quantum systems.
VII Conclusions
For boundary driving, we reveal a universal scaling law for magic, $M_{2}(t)\sim t^{1/z}$, directly tied to the dynamical exponent $z$ of the underlying transport regime and demonstrate that magic serves as a sensitive probe of quantum transport, including the Kardar-Parisi-Zhang universality class.
We introduce a mean-field approach for magic, which isolates the classical contribution to magic through local expectation values. The difference between the full and mean-field magic serves as a direct probe of nonstabilizerness generated by the many-body correlations built into the system.
For bulk dephasing, we demonstrate that dissipation can transiently enhance magic before ultimately driving it toward its steady-state value. Magic exhibits an initial growth, peaks around a characteristic dephasing timescale $t\approx 1/\gamma_{z}$, and then decays exponentially. In systems with conserved magnetization, however, magic is not fully erased and stabilizes at a non-trivial steady-state value. In the Heisenberg limit, we derive an analytical expression capturing this time dependence. Away from the SUSU(2) point, the dynamics are governed by two key timescales: the dephasing time $\tau_{d}\sim 1/\gamma_{z}$, which controls the exponential decay, and the coherence time $\tau_{c}\sim 1/J|\Delta-1|$, which accounts for transient oscillations away from the $SU(2)$ point.
Interestingly, when the system is initialized in configurations such as the Néel state or $(|+\rangle\langle+|)^{\otimes L}$, both belonging to the zero-magnetization sector, the magic exhibits a power-law decay in time. This behavior reflects the gapless character of the Lindbladian dynamics in this regime [see Fig. 4(d,e) and Ref. [40] Sec. IIc]. A more detailed investigation, including the magic dependence on the initial state and system parameters, is the subject of another work.
Acknowledgements.
Appendix A Magic calculation within the vectorized basis